# Cubic* criticality emerging from quantum loop model on triangular lattice

### v2

@@ Abstract
We found in the VP phase both correlation functions decay to a constant value, while exhibit power-law decay at the Cubic* QCP. Interestingly, these two correlators decay with obvious different exponents. For the $t$-term correlation $\langle T(0)T(\mathbf{r})\rangle\sim 1/x^{1+\eta_{T}}$ is consistent with an anomalously large anomalous dimension of the rank-2 tensor of Cubic CFT with $\eta_{T}=1.42$; and for the vison correlation $\langle\bar{v}(0)\bar{v}(\mathbf{r})\rangle\sim 1/x^{1+\eta}$ is consistent with $\eta=0.04$ which is the (2+1)d O(3)/Cubic value of $\eta$ for the order parameter. To access the thermodynamic limit, we depict correlators with different system sizes at the Cubic* QCP in Figs. 2 (c) and 3 (c), and put the small system sizes data of other values of $V$ in SM sup , all these results reveal the $\eta_{T}=1.42$ for $\langle T(0)T(\mathbf{r})\rangle$ and the $\eta=0.04$ for the $\langle\bar{v}(0)\bar{v}(\mathbf{r})\rangle$. On the other hand, inside the QSL phase such as $V=1$ the RK point, both the correlators decay exponentially as shown in Figs. 2 (b) and 3 (b).
@@
Large anomalous dimension means a large scaling dimension as $\Delta_{T}=\frac{1+\eta_{T}}{2}$ for the rank-2 tensor and $\Delta=\frac{1+\eta}{2}$ for the scalar operators of the Cubic/O(3) CFT, our results therefore mean that at the Cubic* QCP, the $t_{1,2,3}$-term is a composite of the fractionalized visons $v_{1,2}$, instead of a well-defined critical mode, and it is the proliferated visons $v$ give rise to the large anomalous dimension of $t$, which serves as a defining signature of the Cubic* transition, different from the conventional Cubic/O(3) QCPs. Similar behavior have been observed in the (2+1)d XY* transition between $\mathbb{Z}_{2}$ QSL and U(1) symmetry-breaking superfluid phase Wang et al. (2017); Isakov et al. (2012); Wang et al. (2018, 2021).
@@
Such fractionalization signature is also vividly seen from the dynamic probes. We measure the dynamic correlation functions $D(\mathbf{k},\tau)$ and $\bar{v}(\mathbf{k},\tau)$ and obtain the dimer and vison spectra via QMC+SAC (details of the scheme is given in SM sup ). Fig. 4 show the obtained spectra across the Cubic* transition. Inside the QSL phase denoted by panels (c) and (f), both spectra exhibit gapped behavior and substantial continua in a large fraction of the momenta along the high-symmetry path. It is interesting to note that the minimal dimer gap is larger than the minimal vison gap due to the fact that dimer is the composite of a pair of visons Feldner et al. (2011); Yan et al. (2021b).
@@
At the Cubic* QCP, the dimer spectra remains gapped as shown in Fig. 4 (b), especially at the $\mathbf{M}_{j=1,2,3}$ points (the seemingly small gap at $\Gamma$ point is due to the finite-size effect), but as shown in Fig. 4 (e), the vison spectra developed a clear gapless mode close to the $\mathbf{M}$ points, as $\mathbf{M}$ points are the ordered wavevector of the VP phase (as explained in Eq. (5)), this critical and gapless vison mode is the dynamic signature of the vison condensation at the Cubic* transition. The contrast between Fig. 4 (b) and (e) explaines the reason why the dimer correlation cannot see the "hidden" VP order and only the vison spectra reveal the translational symmetry breaking of the VP phase. Similar dynamic signature of the $\mathbb{Z}_{2}$ topological order in QSL and the condensation of fractionalized anyons has also been shown in the (2+1)d XY* transition Wang et al. (2017, 2018); Sun et al. (2018); Becker and Wessel (2018); Wang et al. (2021, 2021).
@@
Acknowledgements—We thank Ning Su, Fabien Alet, Rhine Samajdar and Subir Sachdev for insightful discussions on the O($3$) and cubic fixed points and the phase diagram of triangular lattice QLM. XXR, ZY and ZYM acknowledge the support from the Research Grants Council (RGC) of Hong Kong Special Administrative Region of China (Project Nos. 17301721, AoE/P701/20, 17309822, C7037-22GF, 17302223), the ANR/RGC Joint Research Scheme sponsored by RGC of Hong Kong and French National Research Agency (Project No. A_HKU703/22). The K. C. Wong Education Foundation (Grant No. GJTD-2020-01). YQ acknowledges support from the the National Natural Science Foundation of China (Grant Nos. 11874115 and 12174068). JR is supported by Huawei Young Talents Program at IHES. Y.C.W. acknowledges support from Zhejiang Provincial Natural Science Foundation of China (Grant Nos. LZ23A040003). We thank the Beijng PARATERA Tech CO.,Ltd. (URL: https://cloud.paratera.com), the High performance Computing Centre of Beihang Hangzhou Innovation Institute Yuhang, HPC2021 system under the Information Technology Services and the Blackbody HPC system at the Department of Physics, University of Hong Kong for providing computational resources that have contributed to the research results in this paper.
@@ II QMC+SAC SCHEME
[…]
@@
[…]

### v3

@@ Abstract
We found in the VP phase both correlation functions decay to a constant value, while exhibiting power-law decay at the cubic* QCP. Interestingly, these two correlators decay with obvious different exponents. For the $t$-term correlation, $\langle T(0)T(\mathbf{r})\rangle\sim 1/x^{1+\eta_{T}}$ is consistent with an anomalously large anomalous dimension of the rank-2 tensor of cubic CFT with $\eta_{T}=1.42$, and for the vison correlation $\langle\bar{v}(0)\bar{v}(\mathbf{r})\rangle\sim 1/x^{1+\eta}$ is consistent with $\eta=0.04$, which is the (2+1)D O(3)/cubic value of $\eta$ for the order parameter. To access the thermodynamic limit, we depict correlators with different system sizes at the cubic* QCP in Figs. 2 (c) and 3 (c), and put the small system sizes data of other values of $V$ in the SM sup . All these results reveal the $\eta_{T}=1.42$ for $\langle T(0)T(\mathbf{r})\rangle$ and the $\eta=0.04$ for the $\langle\bar{v}(0)\bar{v}(\mathbf{r})\rangle$. On the other hand, inside the QSL phase such as $V=1$, the RK point, both correlators decay exponentially as shown in Figs. 2 (b) and 3 (b).
@@
Large anomalous dimension means a large scaling dimension as $\Delta_{T}=\frac{1+\eta_{T}}{2}$ for the rank-2 tensor and $\Delta=\frac{1+\eta}{2}$ for the scalar operators of the cubic/O(3) CFT, our results therefore mean that at the cubic* QCP, the $t_{1,2,3}$-term is a composite of the fractionalized visons $v_{1,2}$, instead of a well-defined critical mode, and it is the proliferated visons $v$ that give rise to the large anomalous dimension of $t$, which serves as a defining signature of the cubic* transition, different from the conventional cubic/O(3) QCPs. Similar behavior has been observed in the (2+1)D XY* transition between the $\mathbb{Z}_{2}$ QSL and U(1) symmetry-breaking superfluid phase Wang et al. (2017); Isakov et al. (2012); Wang et al. (2018, 2021).
@@
Such a fractionalization signature is also vividly seen from the dynamic probes. We measure the dynamic correlation functions $D(\mathbf{k},\tau)$ and $\bar{v}(\mathbf{k},\tau)$ and obtain the dimer and vison spectra via QMC+SAC (details of the scheme is given in the SM sup ). Figure 4 shows the obtained spectra across the cubic* transition. Inside the QSL phase denoted by Figs. 4 (c) and  4(f), both spectra exhibit gapped behavior and substantial continua in a large fraction of the momenta along the high-symmetry path. It is interesting to note that the minimal dimer gap is larger than the minimal vison gap due to the fact that a dimer is the composite of a pair of visons Feldner et al. (2011); Yan et al. (2021b).
@@
At the cubic* QCP, the dimer spectra remain gapped at the $\mathbf{M}_{j=1,2,3}$ points, as shown in Fig. 4 (b). However, as depicted in Fig. 4 (e), the vison spectra develop a clear gapless mode close to the $\mathbf{M}$ points. Since the $\mathbf{M}$ points are the ordered wave vector of the VP phase [as explained in Eq. (5)], this critical and gapless vison mode serves as the dynamic signature of vison condensation at the cubic* transition. The contrast between Figs. 4 (b) and  4(e) explains why the dimer correlation cannot detect the "hidden" VP order, and only the vison spectra reveal the translational symmetry breaking of the VP phase. Similar dynamic signature of the $\mathbb{Z}_{2}$ topological order in QSL and the condensation of fractionalized anyons have also been demonstrated in the (2+1)D XY* transition Wang et al. (2017, 2018); Sun et al. (2018); Becker and Wessel (2018); Wang et al. (2021, 2021).
@@
Acknowledgements—We thank Ning Su, Fabien Alet, Rhine Samajdar and Subir Sachdev for insightful discussions on the O($3$) and cubic fixed points and the phase diagram of triangular lattice QLM. XXR, ZY and ZYM acknowledge the support from the Research Grants Council (RGC) of Hong Kong Special Administrative Region of China (Project No. 17301721, No. AoE/P701/20, No. 17309822, No. C7037-22GF, 17302223), the ANR/RGC Joint Research Scheme sponsored by RGC of Hong Kong and French National Research Agency (Project No. A_HKU703/22). The K. C. Wong Education Foundation (Grant No. GJTD-2020-01). YQ acknowledges support from the the National Natural Science Foundation of China (Grant No. 11874115 and No. 12174068). JR is supported by Huawei Young Talents Program at IHES. Y.C.W. acknowledges support from Zhejiang Provincial Natural Science Foundation of China (Grant No. LZ23A040003). We thank the Beijng PARATERA Tech CO.,Ltd. (URL: https://cloud.paratera.com), the High performance Computing Centre of Beihang Hangzhou Innovation Institute Yuhang, HPC2021 system under the Information Technology Services and the Blackbody HPC system at the Department of Physics, University of Hong Kong for providing computational resources that have contributed to the research results in this paper.
@@ II QMC+SAC SCHEME
$D(\mathbf{k},\tau)=\int_{0}^{\infty}K(\tau,\omega)S(\mathbf{k},\omega)D\omega,$ (S2)
@@
Note that $D(\mathbf{k},\beta-\tau)=D(\mathbf{k},\tau)$ under this relation as well, thus we only need $\tau$ in the range of $[0,\beta/2]$. For the normalization $D(\mathbf{k},0)=1$, we therefore define $D(\mathbf{k},\omega)=(1+e^{-\beta\omega})S(\mathbf{k},\omega)/\pi$, where $\int_{0}^{\infty}D(\mathbf{k},\omega)D\omega=1$. Then, Eq. S2 is replaced with
$D(\mathbf{k},\tau)=\int_{0}^{\infty}\frac{e^{-\tau\omega}+e^{-(\beta-\tau)\omega}}{1+e^{-\beta\omega}}D(\mathbf{k},\omega)D\omega.$ (S4)